# ../../Bibliographies/main.bib

### v1

@@
Henry Towsner
(Date: April 2, 2025)
@@ 2. Background: Buchholz’s Ordinal
  • $\Omega_{n}<\Omega_{m}$ if $n<m$,
  • $\Omega_{n}<\vartheta_{m}\alpha$ if there is a $\beta\in K_{m}\alpha$ with $\Omega_{n}\leq\beta$,
@@
It will be useful to consider ordinal terms with free variables [^4] This is why we prefer to use ##\##: we could have $\alpha+v=\beta+v$ even when $\alpha\neq\beta$.: we add, for each $n>0$, an additional ordinal term $v_{n}$, understood to have formal cardinality $n$, together with the rule that $\vartheta_{n}\alpha$ cannot contain any $v_{m}$ with $m\geq n$. We extend the ordering by specifying that $v_{n}$ is in $\mathbb{SC}$, is smaller than $\Omega_{n}$, but is larger than any ordinal term with formal cardinality less than $n$; we also set $K_{n}v_{m}=\left\{\begin{array}[]{ll}\{v_{m}\}&\text{if }m<n\\ \emptyset&\text{if }n\leq m\end{array}\right.$, just like $\Omega_{m}$. (Only the first case actually gets applied, because of the restriction on variables appearing in $\vartheta_{n}\alpha$.)
@@
Lemma 2.5 (Key Lemma).
@@
In fact, we need slightly more than this—we need a “relativized” version in which the condition $K_{n}\alpha<\beta$ is replaced by $K_{n}\alpha<\max\{\beta,\delta\}$ for some $\delta$, and corresponding changes are made elsewhere—but this lemma suffices to illustrate the spirit of the argument.
@@
The first part is a straightforward induction on $\alpha$ and $\beta$. Note that $v_{n}$ does not appear inside any subterm $\vartheta_{m}\delta$ with $m\leq n$, so if $\alpha=\vartheta_{m}\alpha^{\prime}$ and $\beta=\vartheta_{m}\beta^{\prime}$, either $m\leq n$, in which case the substitution leaves both ordinal terms unchanged, or $n<m$, in which case $\gamma<\vartheta_{m}\beta^{\prime}$, ensuring the comparison is unchanged.
The second part is essentially the definition.
For the third part, the first part ensures that $\gamma[v\mapsto\vartheta_{n}\alpha]<\beta$, so it suffices to show that $K_{n}(\gamma[v\mapsto\vartheta_{n}\alpha])<\vartheta_{n}\beta$. We have $K_{n}(\gamma[v\mapsto\vartheta_{n}\alpha])\subseteq K_{n}\gamma\cup\{\vartheta_{n}\alpha\}$; $K_{n}\gamma<\vartheta_{n}\beta$ by assumption and $\vartheta_{n}\alpha<\vartheta_{n}\beta$ by the first part. ∎
@@ 3.1. Ordinal Terms
However we need to write terms like the result of the substitution $(\vartheta(\Omega\#v))[v\mapsto\Omega]$; this cannot be $\vartheta(\Omega\#\Omega)$, since the two $\Omega$’s are different. Our analogy to free variables helps us here: we think of $\Omega$ as a free variable, $\vartheta$ as the corresponding binder, and we use de Bruijn indices to distinguish different “levels” of $\Omega$. So we will write $\Omega^{(J)}$, for various natural numbers $J$, to help us distinguish different copies [^6] It feels a trifle silly to trumpet how we’re going to remove the subscripts and then right away replace them with superscripts, but we have actually made a meaningful change here, and the superscripts will behave rather differently. of $\Omega$.
Formally, $J$ counts the number of applications of $\vartheta$ it takes to bind $\Omega$; it is convenient to index this so that it requires $J+1$ applications. We will sometimes use color coding to match applications of $\vartheta$ with the corresponding bound copies of $\Omega$ (and write unbound copies of $\Omega$ in black). So ${\color[rgb]{0,0,1}\vartheta}({\color[rgb]{0,0,1}\Omega}^{(0)}\#v)$ is the function we were considering above, and ${\color[rgb]{0,0,1}\vartheta}({\color[rgb]{0,0,1}\Omega}^{(0)}\#v)[v\mapsto\Omega^{(0)}]={\color[rgb]{0,0,1}\vartheta}({\color[rgb]{0,0,1}\Omega}^{(0)}\#\Omega^{(1)})$. (One confusing aspect of de Bruijn indices, which takes some getting used to, is that the same variable is written differently in different contexts. For instance, in $\Omega^{(0)}\#{\color[rgb]{0,0,1}\vartheta}({\color[rgb]{0,0,1}\Omega}^{(0)}\#\Omega^{(1)})$, the two black copies of $\Omega$ are the same, even though the index has changed: moving inside the $\vartheta$ has incremented the index.)
@@
  • for each natural number $J$, $\Omega^{(J)}$ is in $\mathbb{H}$,
@@
We next need to define the ordering. As a first step, how should we compare $\Omega^{(0)}$ and $\Omega^{(1)}$? In the term ${\color[rgb]{0,0,1}\vartheta}({\color[rgb]{0,0,1}\Omega}^{(0)}\#\Omega^{(1)})$, $\Omega^{(0)}$ is “bound” (that is, has been collapsed by a $\vartheta$) while $\Omega^{(1)}$ is “free”. We always collapse larger cardinals first, so we must have $\Omega^{(0)}>\Omega^{(1)}$. The comparison should depend only on the order of the indices, so we must have $\Omega^{(J)}<\Omega^{(J^{\prime})}$ when $J^{\prime}<J$.
This immediately, of course, makes the full system ill-founded. We will resolve this below by only considering a subclass of ordinal terms.
It is helpful to think of formal cardinality in this system as being a relative notion—$\Omega^{(0)}$ means the current cardinality, $\Omega^{(1)}$ means the cardinality one smaller, and so on [^7] It may be counterintuitive that the order of the indices is flipped like this. We have adopted here the convention which matches de Bruijn indices of variables.. This will come up repeatedly in the definitions below: we will set out to define some notion, but in order to make the induction go through, we will have to keep track of how many times we have gone inside a $\vartheta$.
@@
We define $\mathbb{FC}^{\geq J}(\alpha)$ inductively by:
  • $\mathbb{FC}^{\geq J}(\#\{\alpha_{i}\})=\bigcup_{i}\mathbb{FC}^{\geq J}(\alpha_{i})$,
  • $\mathbb{FC}^{\geq J}(\omega^{\alpha})=\mathbb{FC}^{\geq J}(\alpha)$,
@@
  • $\mathbb{FC}^{\geq J}(\Omega^{(J^{\prime})})=\left\{\begin{array}[]{ll}\emptyset&\text{if }J^{\prime}<J\\ \{J-J^{\prime}\}&\text{if }J\leq J^{\prime}\end{array}\right.$
  • $\mathbb{FC}^{\geq J}(\vartheta\alpha)=\mathbb{FC}^{\geq J+1}\alpha$.
We define $\overline{\mathbb{FC}}^{\geq J}(\alpha)=\sup\mathbb{FC}^{\geq J}(\alpha)$ where $\sup\emptyset=-\infty$.
@@
For instance, $\overline{\mathbb{FC}}(\Omega^{(1)}\#\Omega^{(2)}\#\Omega^{(2)})=-1$—the largest “cardinal” appearing is $\Omega^{(1)}$, so that determines the cardinality, and it’s flipped to cardinality $-1$. We should think of this as saying that this term has cardinality one smaller than the cardinality of our “starting point”. Similarly, $\overline{\mathbb{FC}}({\color[rgb]{0,0,1}\vartheta}({\color[rgb]{0,0,1}\Omega}^{(0)}\#\Omega^{(1)}))=0$. That is, $\mathbb{FC}$ is telling us which uncollapsed cardinals appearing inside $\alpha$, with their levels reinterpreted to be relative to our starting point.
@@
The next thing we need to define is the critical subterms $K\alpha$. Consider the term ${\color[rgb]{1,0,0}\vartheta}{\color[rgb]{0,0,1}\vartheta}({\color[rgb]{0,0,1}\Omega}^{(0)}\#{\color[rgb]{1,0,0}\Omega}^{(1)})$. This is the term we would previous have written as $\vartheta_{1}\vartheta_{2}(\Omega_{2}\#\Omega_{1})$, so it should have no critical subterms. On the other hand, ${\color[rgb]{1,0,0}\vartheta}{\color[rgb]{0,0,1}\vartheta}({\color[rgb]{0,0,1}\Omega}^{(0)})$ is analogous to $\vartheta_{1}\vartheta_{1}\Omega_{1}$, so does have a critical subterm ${\color[rgb]{0,0,1}\vartheta}{\color[rgb]{0,0,1}\Omega}^{(0)}$.
@@
How does $K$ know the difference? The answer is that $K$ knows that it is trying to produce a term $<\Omega^{(0)}$, and so should only pick up subterms which are themselves $<\Omega^{(0)}$. As we define this inductively, our “position” in the relative hierarchy of cardinals changes: say $\vartheta\alpha$ has cardinality too large, so $K\vartheta\alpha$ is supposed to be $K\alpha$; moving inside the $\vartheta$ shifts our position in the cardinals, so the thing we called $\Omega^{(0)}$ outside of $\vartheta\alpha$ is the thing called $\Omega^{(1)}$ inside $\alpha$.
@@
There is a further complication. Consider the term ${\color[rgb]{1,0,0}\vartheta}({\color[rgb]{0,0,1}\vartheta}({\color[rgb]{0,0,1}\Omega}^{(0)}\#\Omega^{(2)})\#{\color[rgb]{1,0,0}\Omega}^{(1)})$. This is analogous [^8] We should highlight here that the notation given by $\mathrm{OT}^{poly}_{\Omega_{\omega}}$ is not trivially isomorphic to $\mathrm{OT}^{poly}_{\Omega_{\omega}}$. For instance, we have no term directly corresponding to $\vartheta_{2}\Omega_{2}$. Our notation here is isomorphic to the subset of the terms from the previous section such that any term $\vartheta_{n}\alpha$ satisfies $n=\max\{\overline{\mathbb{FC}(\alpha)},1\}$. to $\vartheta_{2}(\vartheta_{2}(\Omega_{2}\#\Omega_{1})\#\Omega_{2})$—in particular, we should have $K({\color[rgb]{0,0,1}\vartheta}({\color[rgb]{0,0,1}\Omega}^{(0)}\#\Omega^{(2)})\#{\color[rgb]{1,0,0}\Omega}^{(1)})=\{{\color[rgb]{0,0,1}\vartheta}({\color[rgb]{0,0,1}\Omega}^{(0)}\#\Omega^{(1)})\}$—that is, we ought to shift the index down to reflect that it is no longer inside ${\color[rgb]{1,0,0}\vartheta}$.
@@
So we need to define the notion of shifting the indices in a term up or down to represent “the same term from a different perspective”. [^9] All this shifting might lead one to be skeptical of the entire “relative cardinality” approach. We could instead have tried to define cardinality in an absolute way, so that $\Omega^{(0)}$ means the first cardinality wherever it appears. This would change when shifting has to be done, but would not save us from doing it—for instance, we would have ${\color[rgb]{0,0,1}\vartheta_{0}}({\color[rgb]{0,0,1}\Omega}^{(0)}\#v)[v\mapsto\Omega^{(0)}]={\color[rgb]{0,0,1}\vartheta_{1}}({\color[rgb]{0,0,1}\Omega}^{(1)}\#\Omega^{(0)})$.
@@
We once again need to include a $\cdot^{\geq J}$ restriction—we should only shift unbound cardinals, leaving bound ones alone.
@@
We define $\alpha^{\geq J}_{\pm n}$ by:
  • #$\#\{\alpha_{i}\}^{\geq J}_{\pm n}=\#\{(\alpha_{i})^{\geq J}_{\pm n}\}$,
  • $(\omega^{\alpha})^{\geq J}_{\pm n}=\omega^{\alpha^{\geq J}_{\pm n}}$,
  • $(\Omega^{(J^{\prime})})^{\geq J}_{\pm n}=\left\{\begin{array}[]{ll}J^{\prime}\pm n&\text{if }J^{\prime}\geq J\\ J^{\prime}&\text{if }J^{\prime}<J\end{array}\right.$
  • $(\vartheta\alpha)^{\geq J}_{\pm n}=\vartheta(\alpha^{\geq J+1}_{\pm n})$.
Note that $\alpha^{\geq J}_{-n}$ may not be well-behaved unless $\overline{\mathbb{FC}}^{\geq J}(\alpha)\leq J-n$—for instance, we cannot take $(\vartheta\Omega^{(1)})^{\geq 0}_{-1}$, since reducing the index of $\Omega^{(1)}$ by $1$ would create a collision.
@@
We define $K^{\geq J}\alpha$ inductively by:
  • $K^{\geq J}\#\{\alpha_{i}\}=\bigcup_{i}K^{\geq J}\alpha_{i}$,
  • $K^{\geq J}\omega^{\alpha}=K^{\geq J}\alpha$,
  • $K^{\geq J}\Omega^{(J^{\prime})}=\left\{\begin{array}[]{ll}\{\Omega^{J^{\prime}-(J+1)}\}&\text{if }J^{\prime}\geq J+1\\ \emptyset&\text{if }J^{\prime}<J+1\end{array}\right.$,
  • $K^{\geq J}\vartheta\alpha=\left\{\begin{array}[]{ll}(\vartheta\alpha)^{\geq J}_{-(J+1)}&\text{if }\overline{\mathbb{FC}}^{\geq J+1}(\alpha)<0\\ K^{\geq J+1}\alpha&\text{if }\overline{\mathbb{FC}}^{\geq J+1}(\alpha)\geq 0\end{array}\right.$
The definition of the order is almost unchanged.
@@
  • $\Omega^{(J)}<\Omega^{(J^{\prime})}$ if $J^{\prime}<J$,
  • $\Omega^{(J)}<\vartheta\alpha$ if there is a $\beta\in K^{\geq 0}\alpha$ with $\Omega^{(J)}\leq\beta$,
  • $\vartheta\alpha<\Omega^{(J)}$ if for all $\beta\in K^{\geq 0}\alpha$, $\beta<\Omega^{(J)}$,
@@
    • $\alpha<\beta$ and, for all $\gamma\in K^{\geq 0}\alpha$, $\gamma<\vartheta\beta$, or
    • $\beta<\alpha$ and there is some $\gamma\in K^{\geq 0}\beta$ so that $\vartheta\alpha\leq\gamma$.

### v2

@@
Henry Towsner Department of Mathematics, University of Pennsylvania, 209 South 33rd Street, Philadelphia, PA 19104-6395, USA htowsner@math.upenn.edu http://www.math.upenn.edu/~htowsner
(Date: June 18, 2025)
Partially supported by NSF grant DMS-2054379
@@ 2. Background: Buchholz’s Ordinal
  • $\Omega_{m}<\Omega_{n}$ if $m<n$,
  • $\Omega_{m}<\vartheta_{n}\alpha$ if there is a $\beta\in K_{n}\alpha$ with $\Omega_{m}\leq\beta$,
@@
It will be useful to consider ordinal terms with conventional free variables [^4] This is why we prefer to use ##\##: substitution of variables, at least by ordinals in $\mathbb{H}$, does not alter the normal of an ordinal when we use ##\##.: we add, for each $n>0$, an additional ordinal term $v_{n}$, understood to have formal cardinality $n$, together with the rule that $\vartheta_{n}\alpha$ cannot contain any $v_{m}$ with $m\geq n$. We extend the ordering by specifying that $v_{n}$ is in $\mathbb{SC}$, is smaller than $\Omega_{n}$, but is incomparable to other terms; we also set $K_{n}v_{m}=\left\{\begin{array}[]{ll}\{v_{m}\}&\text{if }m<n\\ \emptyset&\text{if }n\leq m\end{array}\right.$, just like $\Omega_{m}$. (Only the first case actually gets applied, because of the restriction on variables appearing in $\vartheta_{n}\alpha$.)
@@
Lemma 2.5 (Key Lemma, weak version).
@@
In fact, we need slightly more than this—we need a “relativized” version in which the condition $K_{n}\alpha<\beta$ is replaced by $K_{n}\alpha<\max\{\beta,\delta\}$ for some $\delta$, and corresponding changes are made elsewhere.
Definition 2.6.
When $\overline{\mathbb{FC}}(\gamma)<n$, we define $D_{n,\gamma}\beta=\vartheta_{n}(\omega^{\Omega_{n}\#\beta}\#\gamma)$. More generally, for $m\in(0,n)$, we define $D_{m,\gamma}\beta=D_{m,0}D_{m+1,\gamma}\beta$.
When $\overline{\mathbb{FC}}(\gamma)<n$, we define $\alpha\ll^{n}_{\gamma}\beta$ to hold if $\alpha<\beta$ and, for all $m\leq n$, every element of $K_{m}\alpha$ is bounded by $D_{m}\beta$.
Note that this implies that $\ll^{0}_{\gamma}$ is just $<$.
Lemma 2.7 (Key Lemma).
  1. (1)
    If $\alpha<\beta$ and $\delta$ has formal cardinality $<n$ then $\alpha[v_{n}\mapsto\gamma]<\beta[v_{n}\mapsto\gamma]$.
  2. (2)
    If $\overline{\mathbb{FC}}(\delta)<n$, $\alpha\ll^{n}_{\delta}\beta$ are ordinal terms containing no free variables $v_{m}$ with $m\geq n$ then $D_{n,\delta}(\alpha)\ll^{n-1}_{0}D_{n,\delta}(\beta)$.
  3. (3)
    If $\overline{\mathbb{FC}}(\delta)<n$, $\alpha\ll^{n}_{\delta}\beta$ are ordinal terms containing no free variables $v_{m}$ with $m\geq n$, and $\gamma\ll^{n}_{\delta}\beta$ is an oridnal term containing no free variables $v_{m}$ with $m>n$, then $D_{n,D_{n,\delta}(\alpha)}(\gamma[v_{n}\mapsto D_{n,\delta}(\alpha)])\ll^{n-1}_{\delta}D_{n,\gamma}(\beta)$.
@@
  1. (1)
    By a straightforward induction on $\alpha$ and $\beta$. The most difficult case is when $\alpha=\vartheta_{m}\alpha^{\prime}$ and $\beta=\vartheta_{m^{\prime}}\beta^{\prime}$, if $m^{\prime}\leq n$ then $v_{n}$ does not appear $\beta^{\prime}$, so $\beta$ is unchanged after substitution. Also we cannot have $v_{n}$ in $\alpha^{\prime}$ (if we did, we would have $m>n$ and therefore $v_{n}\in K_{n}\alpha^{\prime}$, so $\vartheta_{m^{\prime}}\beta^{\prime}<\vartheta_{m}\alpha^{\prime}$), so $\vartheta_{m}\alpha^{\prime}$ is also unchanged by substitution, so the comparison remains.
    If $m^{\prime}>n$ then the claim follows immediately from the inductive hypothesis.
  2. (2)
    First, observe that $\omega^{\Omega_{n}\#\alpha}\#\delta<\omega^{\Omega_{n}\#\beta}\#\delta$. Also, $K_{n}(\omega^{\Omega_{n}\#\alpha}\#\delta)=K_{n}\alpha\cup K_{n}\beta$, so any element of $K_{n}(\omega^{\Omega_{n}\#\alpha}\#\delta)$ is bounded by an element f either $K_{n}\beta$ or $K_{n}\delta$, and is therefore $<D_{n,\delta}(\beta)$. Therefore $D_{n,\delta}(\alpha)<D_{n,\delta}(\beta)$.
    Finally, if $m<n$ and $\eta\in K_{m}D_{n,\delta}(\alpha)$ then $\eta\in K_{m}\alpha$ or $K_{m}\delta$, so in either case $\eta$ is bounded $D_{m,\delta}(\beta)$.
  3. (3)
    Let $\gamma^{\prime}=\gamma[v_{n}\mapsto D_{n,\delta}(\alpha)]$. We have $\gamma^{\prime}<\beta$, so also $\omega^{\Omega_{n}\#\gamma^{\prime}}\#D_{n,\delta}(\alpha)<\omega^{\Omega_{n}\#\beta}\#\delta$. We have $K_{n}\omega^{\Omega_{n}\#\gamma^{\prime}}\#D_{n,\delta}(\alpha)\subseteq K_{n}\gamma\cup K_{n}\delta\cup K_{n}\alpha$, so in particular every element is bounded by an element of $K_{n}\beta\cup K_{n}\delta$, so $D_{n,D_{n,\delta}(\alpha)}(\gamma^{\prime})<D_{n,\delta}(\beta)$.
    Similarly, if $m<n$ and $\eta\in K_{m}D_{n,D_{n,\delta}}(\gamma^{\prime})$ then $\eta\in K_{m}\gamma\cup K_{m}\delta\cup K_{m}\alpha\cup\{D_{n,\delta}(\alpha)\}$ and is therefore bounded by $D_{m,\delta}(\beta)$.
@@ 3.1. Ordinal Terms
However we need to write terms like the result of the substitution $(\vartheta(\Omega\#v))[v\mapsto\Omega]$; this cannot be $\vartheta(\Omega\#\Omega)$, since the two $\Omega$’s are different. Our analogy to free variables helps us here: we think of $\Omega$ as a free variable, $\vartheta$ as the corresponding binder, and we use de Bruijn indices to distinguish different “levels” of $\Omega$. For reasons that will become apparent, it will be helpful to have our levels go down rather than up, so we will write $\Omega^{(J)}$, for various integers $J\leq 0$, to help us distinguish different copies [^6] It feels a trifle silly to trumpet how we’re going to remove the subscripts and then right away replace them with superscripts, but we have actually made a meaningful change here, and the superscripts will behave rather differently. of $\Omega$.
Formally, $-(J-1)$ counts the number of applications of $\vartheta$ it takes to bind $\Omega$. We will sometimes use color coding to match applications of $\vartheta$ with the corresponding bound copies of $\Omega$ (and write unbound copies of $\Omega$ in black). So ${\color[rgb]{0,0,1}\vartheta}({\color[rgb]{0,0,1}\Omega}^{(0)}\#v)$ is the function we were considering above, and ${\color[rgb]{0,0,1}\vartheta}({\color[rgb]{0,0,1}\Omega}^{(0)}\#v)[v\mapsto\Omega^{(0)}]={\color[rgb]{0,0,1}\vartheta}({\color[rgb]{0,0,1}\Omega}^{(0)}\#\Omega^{(-1)})$. (One confusing aspect of de Bruijn indices, which takes some getting used to, is that the same variable is written differently in different contexts. For instance, in $\Omega^{(0)}\#{\color[rgb]{0,0,1}\vartheta}({\color[rgb]{0,0,1}\Omega}^{(0)}\#\Omega^{(-1)})$, the two black copies of $\Omega$ are the same, even though the index has changed: moving inside the $\vartheta$ has decremented the index.)
@@
  • for each integer $J\leq 0$, $\Omega^{(J)}$ is in $\mathbb{H}$,
@@
We next need to define the ordering. As a first step, how should we compare $\Omega^{(0)}$ and $\Omega^{(-1)}$? In the term ${\color[rgb]{0,0,1}\vartheta}({\color[rgb]{0,0,1}\Omega}^{(0)}\#\Omega^{(-1)})$, $\Omega^{(0)}$ is “bound” (that is, has been collapsed by a $\vartheta$) while $\Omega^{(-1)}$ is “free”. We always collapse larger cardinals first, so we must have $\Omega^{(0)}>\Omega^{(-1)}$. The comparison should depend only on the order of the indices, so we must have $\Omega^{(J)}<\Omega^{(J^{\prime})}$ when $J<J^{\prime}$. (We have chosen to use non-positive indices precisely to make the map $J\mapsto\Omega^{(J)}$ order-preserving rather than order reversing.)
This immediately, of course, makes the full system ill-founded. We will resolve this below by normalizing terms based on the smallest cardinality appearing.
It is helpful to think of formal cardinality in this system as being a relative notion—$\Omega^{(0)}$ means the current cardinality, $\Omega^{(-1)}$ means the cardinality one smaller, and so on. This will come up repeatedly in the definitions below: we will set out to define some notion, but in order to make the induction go through, we will have to keep track of how many times we have gone inside a $\vartheta$.
@@
We define $\mathbb{FC}^{\leq J}(\alpha)$ inductively by:
  • $\mathbb{FC}^{\leq J}(\#\{\alpha_{i}\})=\bigcup_{i}\mathbb{FC}^{\leq J}(\alpha_{i})$,
  • $\mathbb{FC}^{\leq J}(\omega^{\alpha})=\mathbb{FC}^{\leq J}(\alpha)$,
@@
  • $\mathbb{FC}^{\leq J}(\Omega^{(J^{\prime})})=\left\{\begin{array}[]{ll}\{J^{\prime}-J\}&\text{if }J^{\prime}\leq J\\ \emptyset&\text{if }J^{\prime}>J\end{array}\right.$
  • $\mathbb{FC}^{\leq J}(\vartheta\alpha)=\mathbb{FC}^{\leq J-1}\alpha$.
We define $\overline{\mathbb{FC}}^{\leq J}(\alpha)=\sup\mathbb{FC}^{\leq J}(\alpha)$ where $\sup\emptyset=-\infty$.
@@
For instance, $\overline{\mathbb{FC}}(\Omega^{(-1)}\#\Omega^{(-2)}\#\Omega^{(-2)})=-1$—the largest “cardinal” appearing is $\Omega^{(-1)}$. We should think of this as saying that this term has cardinality one smaller than the cardinality of our “starting point”. Similarly, $\overline{\mathbb{FC}}({\color[rgb]{0,0,1}\vartheta}({\color[rgb]{0,0,1}\Omega}^{(0)}\#\Omega^{(-1)}))=0$. That is, $\mathbb{FC}$ is telling us which uncollapsed cardinals appearing inside $\alpha$, with their levels reinterpreted to be relative to our starting point.
@@
The next thing we need to define is the critical subterms $K\alpha$. Consider the term ${\color[rgb]{1,0,0}\vartheta}{\color[rgb]{0,0,1}\vartheta}({\color[rgb]{0,0,1}\Omega}^{(0)}\#{\color[rgb]{1,0,0}\Omega}^{(-1)})$. This is the term we would previous have written as $\vartheta_{1}\vartheta_{2}(\Omega_{2}\#\Omega_{1})$, so it should have no critical subterms. On the other hand, ${\color[rgb]{1,0,0}\vartheta}{\color[rgb]{0,0,1}\vartheta}({\color[rgb]{0,0,1}\Omega}^{(0)})$ is analogous to $\vartheta_{1}\vartheta_{1}\Omega_{1}$, so does have a critical subterm ${\color[rgb]{0,0,1}\vartheta}{\color[rgb]{0,0,1}\Omega}^{(0)}$.
@@
How does $K$ know the difference? The answer is that $K$ knows that it is trying to collect terms $<\Omega^{(0)}$, and so should only pick up subterms which are themselves $<\Omega^{(0)}$. As we define this inductively, our “position” in the relative hierarchy of cardinals changes: say $\vartheta\alpha$ has cardinality too large, so $K\vartheta\alpha$ is supposed to be $K\alpha$; moving inside the $\vartheta$ shifts our position in the cardinals, so the thing we called $\Omega^{(0)}$ outside of $\vartheta\alpha$ is the thing called $\Omega^{(1)}$ inside $\alpha$.
@@
There is a further complication. Consider the term ${\color[rgb]{1,0,0}\vartheta}({\color[rgb]{0,0,1}\vartheta}({\color[rgb]{0,0,1}\Omega}^{(0)}\#\Omega^{(-2)})\#{\color[rgb]{1,0,0}\Omega}^{(-1)})$. This is analogous [^7] We should highlight here that the notation given by $\mathrm{OT}^{poly}_{\Omega_{\omega}}$ is not trivially isomorphic to $\mathrm{OT}_{\Omega_{\omega}}$. to $\vartheta_{2}(\vartheta_{2}(\Omega_{2}\#\Omega_{1})\#\Omega_{2})$—in particular, we should have $K({\color[rgb]{0,0,1}\vartheta}({\color[rgb]{0,0,1}\Omega}^{(0)}\#\Omega^{(-2)})\#{\color[rgb]{1,0,0}\Omega}^{(-1)})=\{{\color[rgb]{0,0,1}\vartheta}({\color[rgb]{0,0,1}\Omega}^{(0)}\#\Omega^{(-1)})\}$—that is, we ought to shift the index down to reflect that it is no longer inside ${\color[rgb]{1,0,0}\vartheta}$.
@@
So we need to define the notion of shifting the indices in a term up or down to represent “the same term from a different perspective”. [^8] All this shifting might lead one to be skeptical of the entire “relative cardinality” approach. We could instead have tried to define cardinality in an absolute way, so that $\Omega^{(0)}$ means the first cardinality wherever it appears. This would change when shifting has to be done, but would not save us from doing it—for instance, we would have ${\color[rgb]{0,0,1}\vartheta_{0}}({\color[rgb]{0,0,1}\Omega}^{(0)}\#v)[v\mapsto\Omega^{(0)}]={\color[rgb]{0,0,1}\vartheta_{1}}({\color[rgb]{0,0,1}\Omega}^{(-1)}\#\Omega^{(0)})$.
@@
We once again need to include a $\cdot^{\leq J}$ restriction—we should only shift unbound cardinals, leaving bound ones alone.
@@
We define $\alpha^{\leq J}_{\pm n}$ by:
  • #$\#\{\alpha_{i}\}^{\leq J}_{\pm n}=\#\{(\alpha_{i})^{\leq J}_{\pm n}\}$,
  • $(\omega^{\alpha})^{\leq J}_{\pm n}=\omega^{\alpha^{\leq J}_{\pm n}}$,
  • $(\Omega^{(J^{\prime})})^{\leq J}_{\pm n}=\left\{\begin{array}[]{ll}\Omega^{(J^{\prime}\pm n)}&\text{if }J^{\prime}\leq J\\ \Omega^{(J^{\prime})}&\text{if }J^{\prime}>J\end{array}\right.$
  • $(\vartheta\alpha)^{\leq J}_{\pm n}=\vartheta(\alpha^{\leq J-1}_{\pm n})$.
Note that $\alpha^{\leq J}_{+n}$ may not be well-behaved unless $\overline{\mathbb{FC}}^{\leq J}(\alpha)\leq-n$—for instance, we cannot take $(\vartheta\Omega^{(-1)})^{\leq 0}_{+1}$, since increasing the index of $\Omega^{(-1)}$ by $1$ would create a collision. (Maybe more directly, we cannot take $(\Omega^{(0)})^{\leq 0}_{+1}$ because $\Omega^{(1)}$ is not a term.)
@@
We define $K^{<J}\alpha$ inductively by:
  • $K^{<J}\#\{\alpha_{i}\}=\bigcup_{i}K^{<J}\alpha_{i}$,
  • $K^{<J}\omega^{\alpha}=K^{<J}\alpha$,
  • $K^{<J}\Omega^{(J^{\prime})}=\left\{\begin{array}[]{ll}\{\Omega^{(J^{\prime}-(J-1))}\}&\text{if }J^{\prime}<J\\ \emptyset&\text{if }J^{\prime}\geq J\end{array}\right.$,
  • $K^{<J}\vartheta\alpha=\left\{\begin{array}[]{ll}(\vartheta\alpha)^{\leq 0}_{+(1-J)}&\text{if }\overline{\mathbb{FC}}^{\leq 0}(\vartheta\alpha)<J\\ K^{<J-1}\alpha&\text{if }\overline{\mathbb{FC}}^{\leq 0}(\vartheta\alpha)\geq J\end{array}\right.$
To help motivate the numbers in the definition above, keep in mind that we are interested in using $K^{<0}\alpha$ to help us understand $\vartheta\alpha$. $K^{<J}\beta$ means we are looking at some subterm of $\alpha$ which is inside $-J$ applications of $\vartheta$, and when we pull subterms into the set $K^{<J}\beta$, we wish to shift them by enough so that we can correctly compare them to the term $\vartheta\alpha$.
Having adjusted the definitions, the definition of the order is almost unchanged.
@@
  • $\Omega^{(J)}<\Omega^{(J^{\prime})}$ if $J<J^{\prime}$,
  • $\Omega^{(J)}<\vartheta\alpha$ if there is a $\beta\in K^{<0}\alpha$ with $\Omega^{(J)}\leq\beta$,
  • $\vartheta\alpha<\Omega^{(J)}$ if for all $\beta\in K^{<0}\alpha$, $\beta<\Omega^{(J)}$,
@@
    • $\alpha<\beta$ and, for all $\gamma\in K^{<0}\alpha$, $\gamma<\vartheta\beta$, or
    • $\beta<\alpha$ and there is some $\gamma\in K^{<0}\beta$ so that $\vartheta\alpha\leq\gamma$.